OpenStudy (anonymous):

Find the inverse of the one to one function. f(x)=6x-8

jimthompson5910 (jim_thompson5910):

\[\Large f(x)=6x-8\] \[\Large y=6x-8\] \[\Large x=6y-8\] \[\Large x+8=6y\] \[\Large \frac{x+8}{6}=y\] \[\Large y=\frac{x+8}{6}\] \[\Large f^{-1}(x)=\frac{x+8}{6}\]